The measure of the supplement of an angle is 40 more than two times the measure of the complement of an angle. Find the measure of the angle.
Answer by Flannery(124)

Let C stand for compliment.
Let S stand for suppliment.
Let A stand for angle.
A=90-C
S=A+2C+40
First solve for compliment.
180=90-C+2C+40
140=90-C+2C
140=90+C
50=C
90 degrees minus the compliment is the angle's measure.
90 minus 50 is 40.
40 degrees is the angle's measure.
Check by solving for suppliment.
S=2C+40
S=100+40
S=140
180 degrees minus the suppliment is the angle's measure.
180 minus 140 equals 40.
Fourty is the angle's measure.
Because 40+50=90 AND 40+2(50)+40=180
A=40
C=50
S=140
Q.E.D